Tyrion:

Solve the equation the square root of the quantity x plus 3 plus 4 equals 5 for the variable. Show each step of your solution process.

Tyrion:

is dis it?? √(x + 3) = 5 - 4 √(x + 3) = 1 (√(x + 3))2 = (1)2 x + 3 = 1 x = 1 - 3 x = -2